OpenStudy (anonymous):

Solve negative two-thirds times a plus two-fifths equals three-fourths for a.

OpenStudy (anonymous):

I Need This Answer Please I Need This Test Answer So I Can Pass This Courses

OpenStudy (anonymous):

Is Anyone Going To Help Me!!??

OpenStudy (anonymous):

Well If You Do Help Me I Cannot Give U Anything But Please Have A Heart I Cannt Fail These Classes

OpenStudy (mertsj):

\[-\frac{2}{3}a+\frac{2}{5}=\frac{3}{4}a\]

OpenStudy (anonymous):

Like Help Asap ! Im In Thaa Middle Of A Test ! Wait Whats The Answer???

OpenStudy (mertsj):

No one on OS is going to help you cheat.

OpenStudy (anonymous):

Oh ! Mi Gosh Whatever !

OpenStudy (mertsj):

If you cannot fail this class, you should have learned the material as you went through the class.

OpenStudy (anonymous):

I DO Study The Material Its Sooo Hard ! Whatever Nobody Asked You Okay I Will Ask Another Website ! Gudbye!!!!!!!!!!!!!!!!!!

OpenStudy (mertsj):

I will tell you the steps but you must do them yourself: 1. add 2/3a to both sides 2. Multiply both sides by the reciprocal of the coefficient of a

OpenStudy (anonymous):

Okay Thank You So MuCH!